# Fast pseudothermalization

### v1

@@ III Thermalization of subsets
The probability $p_{n}$ of a $n\times m$ random matrix $M$ over $\mathbb{F}_{2}$ with $m>n$ having the rank $n$ is given by
$p_{n}=\prod_{j=1}^{n}\left(1-q^{m-j+1}\right),$ (3)
where $q$ is the probability of each element of the matrix having zero.
@@
Let us construct a full rank matrix $M$ by sequentially sampling from the first row to the last row. Since $M$ is full rank, its rows are all linearly independent. Let us assume that we have sampled the first $l$ rows which are all linearly independent. Without loss of generality, we can reorder the columns of $M$ so that the first $l\times l$ sub-matrix of $M$ is a full rank matrix. Then, the probability $p_{l+1}$ of sampling a row vector that is linearly independent from the previously sampled rows is given by
$p_{l+1}=1-q^{m-l},$ (4)
since we only need to avoid sampling a row with vanishing elements at $[l+1,m]$. Therefore, $p_{n}$ is given by
$p_{n}=\prod_{l=1}^{n}p_{l}=\prod_{l=1}^{n}\left(1-q^{m-l+1}\right).$ (5)
@@
Algorithm 2 performs random bit-flipping of $t$ distinct copies of $\{0,1\}^{k}\times 0^{n-k}$ in two stages. In the first stage, it takes the first $k$ bits as control bits and the remaining $n-k$ bits as target bits for the MCX gates. For $t$ randomly chosen distinct copies of $\{0,1\}^{k}\times 0^{n-k}$, the first $k$ bits have random binary values for $k\gg n$, causing the MCX gates to randomly flip and thermalize the bits in positions $[k+1,n]$. In the second stage, the algorithm reverses the role of the bits, using the thermalized bits in $[k+1,n]$ as control bits and the first $k$ bits as target bits, thereby thermalizing the first $k$ bits.
@@
For $k=\omega(\log n)$ and $t\leq k/2$, Algorithm 2 with $\alpha t=\omega(\log n)\leq k/2$, and $m=2$ thermalizes all bits of $t$ distinct random copies of $\{0,1\}^{k}\times 0^{n-k}$ with a failure probability negligible in $n$ and $\omega(n\log n)$ CCX gates.
@@
Since $2^{k}$ is super-polynomially greater than $t$, and $t$ copies are sampled from $\{0,1\}^{k}\times 0^{n-k}$, their first $k$ bits are random. Thus, $X$ is a random matrix over $\mathbb{F}_{2}$ whose entries are independent and non-trivial with the probability of $1/4$. Due to Lemma 1, the probability $p_{t}$ of such a matrix having the full rank is given by
$p_{t}=\prod_{l=1}^{t}\left(1-\left(\frac{3}{4}\right)^{\alpha t-l+1}\right).$ (6)
Since we have $\alpha t=\omega(\log n)$, for sufficiently large $n$, this can be approximated by
$\begin{split}p_{n}&\approx\exp\left(-\sum_{l=1}^{t}\left(\frac{3}{4}\right)^{\alpha t-l+1}\right)\\ &=\exp\left(-4\left(\frac{3}{4}\right)^{\alpha t}\left(1-\left(\frac{3}{4}\right)^{t}\right)\right)\\ &\approx 1-\operatorname{negl}(n).\end{split}$ (7)
@@
Let $\{\textbf{x}_{l}\}_{l=1}^{t}$ be $t$ distinct copies of $\{0,1\}^{k}\times 0^{n-k}$. Let us execute $\operatorname{RMC}(n,1,k,m)$ function $\alpha t$ times and get $\{C_{l}\}_{l=1}^{\alpha t}$ and $\{T_{l}\}_{l=1}^{\alpha t}$. Similar to the proof of Theorem 1, we can construct a $t\times\alpha t$ matrix $X$ over $\mathbb{F}_{2}$ whose element $(p,q)$ is one if $((\textbf{x}_{p})_{i_{1,q}},\cdots,(\textbf{x}_{p})_{i_{m,q}})=(x_{i_{1,q}},\cdots,i_{i_{m,q}})$ and zero otherwise. Each element of $X$ is one with the probability higher than $1/t$. Thus, the probability $p_{t}$ of $X$ having the full rank is lower bounded by
$p_{t}\geq\prod_{l=1}^{t}\left(1-\left(1-\frac{1}{t}\right)^{\alpha t-l+1}\right).$ (8)
For a sufficiently large $n$, this can be approximated by
$\begin{split}p_{n}&\gtrsim\exp\left(-\sum_{l=1}^{t}e^{-\alpha+\frac{l-1}{t}}\right)\\ &=\exp\left(-e^{-\alpha}\frac{e-1}{e^{1/t}-1}\right)\\ &=1-\operatorname{negl}(n).\end{split}$ (9)
Thus, $X$ is full rank with a failure probability negligible in $n$, and all bits in $[k+1,n]$ of the $t$ copies are thermalized by applying $\alpha t$ MCX gates with $m$ conditional bits according to $\{C_{l}\}_{l=1}^{\alpha t}$ and $\{T_{l}\}_{l=1}^{\alpha t}$. The remaining bits in $[1,k]$ can be thermalized by applying MCX gates based on results of $\alpha t$ runs of $\operatorname{RMC}(n,k+1,n,m)$. ∎
@@ Acknowledgements.
W.L. and G.Y.C. are supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2002-05 and SSTF-BA2401-03, the NRF of Korea (Grants No. RS-2023-00208291, No. 2023M3K5A1094810, No. 2023M3K5A1094813, No. RS-2024-00410027, No. RS-2024-00444725) funded by the Korean Government (MSIT), the Air Force Office of Scientific Research under Award No. FA2386-22-1-4061, and Institute of Basic Science under project code IBS-R014-D1. H.K. is supported by the KIAS Individual Grant No. CG085301 at Korea Institute for Advanced Study and National Research Foundation of Korea (Grants No. 2023M3K5A109480511 and No. 2023M3K5A1094813).

### v2

@@ III Thermalization of subsets
The probability $p_{l}$ of a $l\times m$ random matrix $M$ over $\mathbb{F}_{2}$ with $m\gg l$ and $l=\operatorname{poly}(n)$ for some $n$ having the full rank is lower bounded by
$p_{l}\leq\exp\left(-\frac{q^{m+1}(q^{-l}-1)}{p}-\frac{sp^{\tilde{q}m+1}q^{\tilde{p}m-1}}{(1-s)^{2}}\right).$ (3)
with a failure probability negligible in $n$, with the non-trivial element probability $p\leq\frac{1}{4}$, $q=1-p$, $\tilde{p}=(1+\epsilon)p$, $\tilde{q}=1-\tilde{p}$, $s=(pq)^{\tilde{q}}$, and $mp=\omega(\log n)$ for some $10<\epsilon<$ 1.
@@
Let us construct a full rank matrix $M$ by sequentially sampling from the first row to the last row. Since $M$ is full rank, its rows are all linearly independent. Let us assume that we have sampled the first $r$ rows which are all linearly independent. Without loss of generality, we can reorder the columns of $M$ so that the first $r\times r$ sub-matrix of $M$ is a full rank matrix. Then, for $m\gg l$, the probability $p_{r+1}$ of sampling a row vector that is linearly independent from the previously sampled rows is given by
$p_{r+1}=1-q^{m-l}-\sum_{r^{\prime}=1}^{l}p_{r^{\prime},\mathrm{ind}}.$ (4)
The second term in the right hand side is the probability of sampling a row with vanishing elements at $[r+1,m]$. $p_{r^{\prime},\mathrm{ind}}$ is the probability of not sampling the previously sampled $r^{\prime}$-th row. The number $N_{r^{\prime}}$ of ones of the $r^{\prime}$-th row is smaller than $(1+\epsilon)np$ for some small $\epsilon>0$ with a failure probability negligible in $n$ due to the Chernoff bound of
$\operatorname{Pr}\left[N_{r^{\prime}}>(1+\epsilon)mp\right]\leq\exp(-\frac{\epsilon^{2}mp}{2+\epsilon}).$ (5)
Let $\tilde{p}$ and $\tilde{q}$ be $(1+\epsilon)p$ and $1-\tilde{p}$, respectively. Then, we have
$p_{r^{\prime},\mathrm{ind}}\geq p^{\tilde{q}(m-r)}q^{\tilde{p}(m-r)}\times pq^{r-1}$ (6)
with a probability higher than $1-\operatorname{negl}(n)$. This directly implies
$p_{r+1}\geq 1-q^{m-r}-lp^{\tilde{q}(m-r)+1}q^{\tilde{p}m+\tilde{q}r-1}.$ (7)
as well as
$\begin{split}p_{l}&\geq\prod_{r=0}^{l-1}\left(1-q^{m-r}-rp^{\tilde{q}(m-r)+1}q^{\tilde{p}m+\tilde{q}r-1}\right)\\ &\approx\exp\left(-\sum_{r=0}^{l-1}\left(q^{m-r}+rp^{\tilde{q}(m-r)+1}q^{\tilde{p}m-\tilde{q}r-1}\right)\right)\\ &=\exp\left(-\frac{1}{p}q^{m+1}(q^{-l}-1)\right)\\ &\quad\times\exp\left(-p^{\tilde{q}m+1}q^{\tilde{p}m-1}\frac{s(1-ls^{l-1}(1-s))}{(1-s)^{2}}\right)\end{split}$ (8)
with a probability higher than $1-\operatorname{negl}(n)$. For $p\leq 1/4$ and large $l$, this becomes
$p_{l}\geq\exp\left(-\frac{q^{m+1}(q^{-l}-1)}{p}-\frac{sp^{\tilde{q}m+1}q^{\tilde{p}m-1}}{(1-s)^{2}}\right).$ (9)
@@
Algorithm 2 performs random bit-flipping of $t$ distinct copies of $\{0,1\}^{k}\times 0^{n-k}$ in two stages. In the first stage, it takes the first $k$ bits as control bits and the remaining $n-k$ bits as target bits for the MCX gates. For $t$ randomly chosen distinct copies of $\{0,1\}^{k}\times 0^{n-k}$, the first $k$ bits have random binary values for $k\ll n$, causing the MCX gates to randomly flip and thermalize the bits in positions $[k+1,n]$. In the second stage, the algorithm reverses the role of the bits, using the thermalized bits in $[k+1,n]$ as control bits and the first $k$ bits as target bits, thereby thermalizing the first $k$ bits.
@@
For $k=\omega(\log n)$ and $t\leq k/2$, Algorithm 2 with $\alpha t=\omega(\log n)\leq k/2$, $\alpha\gg 1$, and $m=2$ thermalizes all bits of $t$ distinct random copies of $\{0,1\}^{k}\times 0^{n-k}$ with a failure probability negligible in $n$ and $\omega(n\log n)$ CCX gates.
@@
Since $2^{k}$ is super-polynomially greater than $t$, and $t$ copies are sampled from $\{0,1\}^{k}\times 0^{n-k}$, their first $k$ bits are random. Thus, $X$ is a random matrix over $\mathbb{F}_{2}$ whose entries are independent and non-trivial with the probability of $1/4$. Due to Lemma 1, the probability $p_{t}$ of such a matrix having the full rank is lower bounded by Eq. (1) with $p=1/4$, $l=t$, $m=\alpha t$, and $10<\epsilon<$ 1,
$\begin{split}p_{t}&\geq\exp\left(-4\left(\frac{3}{4}\right)^{\alpha t+1}\left(\left(\frac{3}{4}\right)^{-t}-1\right)\right)\\ &\quad\times\exp\left(-\frac{s}{(1-s)^{2}}\frac{3^{(1+\epsilon)\alpha/4t-1}}{4^{\alpha t}}\right)\end{split}$ (10)
with a probability higher than $1-\operatorname{negl}(n)$. Since we have $\alpha t=\omega(\log n)$, for sufficiently large $n$, $p_{t}$ is given by
$p_{t}=1-\operatorname{negl}(n).$ (11)
@@
Let $\{\textbf{x}_{l}\}_{l=1}^{t}$ be $t$ distinct copies of $\{0,1\}^{k}\times 0^{n-k}$. Let us execute $\operatorname{RMC}(n,1,k,m)$ function $\alpha t$ times and get $\{C_{l}\}_{l=1}^{\alpha t}$ and $\{T_{l}\}_{l=1}^{\alpha t}$. Similar to the proof of Theorem 1, we can construct a $t\times\alpha t$ matrix $X$ over $\mathbb{F}_{2}$ whose element $(p,q)$ is one if $((\textbf{x}_{p})_{i_{1,q}},\cdots,(\textbf{x}_{p})_{i_{m,q}})=(x_{i_{1,q}},\cdots,i_{i_{m,q}})$ and zero otherwise. Each element of $X$ is one with the probability higher than $1/t$. Thus, for large enough $t$, the full rank probability of $X$ in Eq. (3) with $l=t$, $m=\alpha t$, and $10<\epsilon<$ 1 is lower bounded by
$\begin{split}p_{t}&\geq\exp\left(-t\left(1-\frac{1}{t}\right)^{\alpha t+1}\left(\left(1-\frac{1}{t}\right)^{-t}-1\right)\right)\\ &\quad\times\exp\left(-\frac{s(1/t)^{(t-1-\epsilon)\alpha+1}(1-1/t)^{(1+\epsilon)\alpha-1}}{(1-s)^{2}}\right)\\ &\approx\exp\left(-(t-1)e^{-\alpha}\left(e-1\right)\right)\\ &\quad\times\exp\left(-\frac{(t-1)^{(1+\epsilon)\alpha-1}}{t^{\alpha t}}\frac{s}{(1-s)^{2}}\right)\\ &=1-\operatorname{negl}(n)\end{split}$ (12)
with a probability higher than $1-\operatorname{negl}(n)$. Thus, $X$ is full rank with a failure probability negligible in $n$, and all bits in $[k+1,n]$ of the $t$ copies are thermalized by applying $\alpha t$ MCX gates with $m$ conditional bits according to $\{C_{l}\}_{l=1}^{\alpha t}$ and $\{T_{l}\}_{l=1}^{\alpha t}$. The remaining bits in $[1,k]$ can be thermalized by applying MCX gates based on results of $\alpha t$ runs of $\operatorname{RMC}(n,k+1,n,m)$. ∎
@@ Acknowledgements.
We thank Byungmin Kang for helpful discussions. W.L. and G.Y.C. are supported by Samsung Science and Technology Foundation under Project Number SSTF-BA2002-05 and SSTF-BA2401-03, the NRF of Korea (Grants No. RS-2023-00208291, No. 2023M3K5A1094810, No. 2023M3K5A1094813, No. RS-2024-00410027, No. RS-2024-00444725) funded by the Korean Government (MSIT), the Air Force Office of Scientific Research under Award No. FA2386-22-1-4061, and Institute of Basic Science under project code IBS-R014-D1. H.K. is supported by the KIAS Individual Grant No. CG085301 at Korea Institute for Advanced Study and National Research Foundation of Korea (Grants No. 2023M3K5A109480511 and No. 2023M3K5A1094813).